OpenStudy (anonymous):

Solve using the square root property. 6x^2 + 4 = 0. I got 6x = 2i. Is this correct?

OpenStudy (tkhunny):

Well, since it isn't ataully solved, that would have to be a "no". Never give up until you get "x = ". "6x = " is no good. \(6x^{2} + 4 = 0\) Factor 2 \(2\cdot(3x^{2} + 2) = 0\) Factor as Difference of Squares \(2\cdot(\sqrt{3}x + \sqrt{2}i)\cdot(\sqrt{3}x - \sqrt{2}i) = 0\) Now what?

zepdrix (zepdrix):

Swan, I'm assuming you did this,\[\large 6x^2+4=0\]Subtracting 4 from each side gives us,\[\large 6x^2=-4\]Then you took the square root of both sides,\[\huge \color{red}{6}x=\pm 2i\]You forgot to take the square root of the 6.

OpenStudy (mathstudent55):